Two narrow slits 55 μm apart are illuminated with light of wavelength 505 nm .

Part A:

What is the angle of the m = 2 bright fringe in radians?

Part B:

What is the angle of the m = 2 bright fringe in degrees?
Answer #1

for double slit interferrence,

constructive interferrence (which results in bright fringes) occur when following condition is satisfied.

d*sin(theta)=m*lambda

where d=distance between the two slits=55 um=55*10^(-6) m

theta=angle of the m th order fringe

m=order of the fringe

lambda=wavelength

using the repsective values provided as per the question,

55*10^(-6)*sin(theta)=2*505*10^(-9)

==>sin(theta)=0.018363

part a:

sin(theta)=0.018363

==>theta=0.018364 radians

part b:

sin(theta)=0.018363

==>theta=1.0522 degrees